OpenStudy (anonymous):

find the x and y intercepts and graph the equation plotting the intercepts -4x+5y=-20

OpenStudy (anonymous):

x intercept at (5,0) and y intercept at (0,-4).

OpenStudy (anonymous):

Rearrange the equation to make y the subject... 5y = 4x - 20 y = (4/5)x - 4 When x = 0, what does y equal? y = (4/5)(0) - 4 = -4... (0,-4) Rearrange the equation to make x the subject... 4x = 5y + 20 x = (5/4)y + 5 when y = 0, what does x equal? x = (5/4)(0) + 5 = 5... (5, 0)

OpenStudy (anonymous):

let x=0 solve y, and then let y=0 solve x